Grade 12Electric Current

If the length of the conductor becomes double calculate the value of the new resistance.

Eknoor singh
8 Years agoGrade 12

2 Answers

Athul Joseph
8 Years ago
R-reaistance,ฯ-resistivityVolume=const.A*L=A`*L/2A`=2AInitial resistance,R=ฯL/AFinal resistance,R`=ฯL/4AR`=R/4.Hope u understood.
Khimraj
8 Years ago
Rsistance of a wire is directly proportional to the legth of wire.
If length is doubled by adding same wire then new reistance will be doubled.
and if length is doubled by streching then resistance will be four times of initial.